OpenStudy (anonymous):

U(x,y)=xy^2 MUx=y^2, MUy=2xy Price per unit of good A=2 and for B=4. Income is 1200. How to calculate optimal quantity of good A and good B for this consumer?

OpenStudy (anonymous):

MU of x / MU of y = Y/ 2x this should be equal to price ratio P of x / P of y = 1 / 2